Use cylindrical coordinates to calculate ∫∫∫W f(x,y,z) dV for the given function and region:

f(x,y,z)=z, x^2+y^2≤z≤16
1
Expert's answer
2020-11-17T06:07:34-0500

projection of the region onto the oxy plane:



let's move on to cylindrical coordinates:

"\begin{array}{l}
x = r\cos \varphi ;\,\,y = r\sin \varphi ; z=z\\
0 \le \varphi \le 2\pi ,\,\,0 \le r \le 4,\,\,{x^2} + {y^2} \le z \le 16 \Rightarrow {r^2} \le z \le 16
\end{array}"

Then

"\begin{array}{l}
\mathop {\int {\int {\int {zdV} } } }\limits_V = \int\limits_0^{2\pi } {d\varphi } \int\limits_0^4 {rdr} \int\limits_{{r^2}}^{16} {zdz} = \frac{1}{2}\int\limits_0^{2\pi } {d\varphi } \int\limits_0^4 {r\left. {{z^2}} \right|_{{r^2}}^{16}dr} = \frac{1}{2}\left. \varphi \right|_0^{2\pi } \cdot \int\limits_0^4 {r({{16}^2} - {r^4})dr} = \\
 = \frac{{2\pi }}{2}\int\limits_0^4 {\left( {256r - {r^5}} \right)} dr = \pi \left( {128\left. {{r^2}} \right|_0^4 - \left. {\frac{{{r^6}}}{6}} \right|_0^4} \right) = \pi \left( {128 \cdot 16 - \frac{{4096}}{6}} \right) = \frac{{4096\pi }}{3}
\end{array}" Answer: "\frac{{4096\pi }}{3}"